OpenStudy (anonymous):

True or False? The value pi/4 is a solution for the equation 3√2 cos θ +2=-1. ***My Answer: False*** Is this correct?

OpenStudy (p0sitr0n):

well, you basically need to check the equality, so replacing gives cos(pi/4)=sqrt(2)/2 then 3root(2)*root(2)/2+2=3*1+2=5 and not 1.

OpenStudy (p0sitr0n):

so you are correct

OpenStudy (anonymous):

I see now! Thank you for the help!